If f(x)=|x-2| & g(x)=f[f(x)] then for 2
A)1
B)-1
C)0
D)none of these
 

Grade:12

1 Answers

seetaram dantu
7 years ago
f(x) = | x-2|  , f(2) =|2-2|= 0  then g(x) = f(f(2) ==> g(x) = f[0]
we know that f(x)=|x-2| ==>f(0) = |0-2|= 2
 
therefore, g(x)= 2 and  answer is none of these